Helping my daughter and need your help. Need help solving this problem. I need to solve each system using the elimination method. Write no solution or infinitely many solutions where appropriate.
1. 2x + 5y = -1
x + 2y = 0
Thanks

Answer by Nate(3500)

To use the elimination method, you need to get an opposite of a coefficient with the same coefficient.
x + 2y = 0
2x + 4y = 0
-2x - 4y = 0
Now, combine that with the other equation:
2x + 5y = -1
-2x - 4y = 0
0x + 1y = -1
y = -1
Plug:
x + 2y = 0
x - 2 = 0
x = 2